# Cooper-pair splitters as circuit elements for realizing topological superconductors

### v1

@@ Abstract
Advances in materials and fabrication of superconducting devices allows the exploration of novel quantum effects in synthetic superconducting systems beyond conventional Josephson junction arrays. As an example, we introduce a new circuit element, the Y-splitter, a superconducting loop with three leads and three Josephson junctions, smaller or comparable in size to the superconducting coherence length of the material. By tuning magnetic flux through an array of Y-splitters, Cooper-pair transport can be made to interfere destructively, while spatially separated split Cooper pairs propagate coherently. We consider an array of Y-splitters connected in a two-dimensional star [Archimedean (3,$12^{2}$)] geometry, deformable into the kagome lattice, and find a rich phase diagram that includes topological superconducting phases with Chern numbers ±$\pm 2$. Experimental realization appears feasible.
Introduction: The suppression of dissipation in superconductors is key to building electrical circuits that behave quantum mechanically, making these circuits leading contenders for quantum computing architectures [1, 2, 3, 4]. To date, the collection of non-dissipative superconducting circuit elements includes capacitors, inductors, and Josephson junctions. In describing these circuits theoretically, the underlying pairing of electrons is concealed; aside from occasional factors of two in formulas, electron pairing is assumed then forgotten, except insofar as the separation of pairs is regarded as an undesirable source of decoherence and parity non-conservation.
@@
Figure 1: The Y-splitter, the network component introduced in this work, in two sizes, small (a,c) and large (b,d) compared to the coherence length, $\xi$, and two Josephson coupling strengths, strong (a,b) and weak (c,d). Blue circle represents the coherence length, red segments the insulating barrier thickness. Dominant contribution of Cooper pair transport from the left to the upper arm shown as orange trajectories. The flux $\phi$ through the triangular loop controls interference of single electrons and Cooper pairs. Applying 1/2 (mod 1) flux quanta through the triangular loop leads to destructive interference of Cooper pairs but not single electrons.
@@
The single Y-splitter: The properties of the individual Y-splitter depends on the size of the loop relative to the superconducting coherence length $\xi$ as well as the Josephson coupling strength of its three junctions (assumed to be equal). Four cases are represented in Fig. 1, small (a, c) versus large (b, d) loops, and small (c, d) versus large (a, b) Josephson couplings. In case (a), where the loop is small compared to $\xi$ and couplings are strong (indicated by thin red barriers), Cooper pairs split then recombine while retaining phase coherence, with the phase difference of the separated electrons, accumulated at the junctions, controlled by flux, $\phi$, through the loop, periodic in the single-electron flux quantum, $hc/e$. Cooper pairs can also travel together along each arm, obeying the usual Josephson relations at the junctions within each arm. In case (b), with a large loop compared to $\xi$, coherent pair splitting and recombining is exponentially suppressed. In case (c), coherent splitting is allowed by the small loop but is suppressed compared to pair transport by the weak junctions [thicker red junctions in (c,d)]. This is because transport along a single edge is a second order tunneling process, while transferring two electrons along different paths is a third order process. For case (d), the large loop and weak tunneling doubly suppress coherent splitting.
The Y-splitter network: Next, we construct a hexagonal 2D array of Y-splitters, which is equivalent to an Archimedean (3,$12^{2}$) or ”star” lattice of wires, as depicted in Fig. 2(a). Notice that as the length of the segment connecting neighboring junctions is shrunk to zero and the two connected Y-junctions become an X-shaped wire crossing, the network morphs into a kagome lattice shown in Fig. 2(b). The kagome geometry to topological superconductivity (see, e.g., Ref. [35]) in tight-binding fermionic models; by breaking the Cooper pairs with the Y-splitters, we access the fermionic degrees of freedom in an artificially built lattice.
In the kagome configuration [Fig. 2(b)], the unit cell area $A_{\rm cell}=A_{\triangleright}+A_{\triangleleft}+A_{\varhexagon}$ comprises two triangles of area $A_{\triangleright}=A_{\triangleleft}$ each and an hexagon of area $A_{\varhexagon}=6\,A_{\triangleright}$. We consider applied perpendicular magnetic fields such that the flux per unit cell is an integer multiple of the flux quantum $\Phi_{0}=\frac{hc}{e}$, a point to which we return when discussing the experimental considerations for such systems. In this situation, we can study model Hamiltonians that do not demand the use of magnetic super cells (in contrast to Hofstadter-like models. This case of zero net flux (mod one flux quantum) per unit cell is equivalent to flux $\phi$ per triangle and $-2\phi$ per hexagon. (We henceforth work in units in which the flux $\phi$ is an angular variable, $\phi=2\pi\Phi/\Phi_{0}$, where $\Phi$ is the physical flux through the triangles, pointing into the page.)
@@
Figure 3: Phase diagram of the effective model. The diagram shows the chirality $\chi$ and Chern numbers $\cal{C}$ as functions of the flux $\phi$ and ratio $\Gamma/\Delta$. The chirality takes specific values $\chi_{32}=0,\pm 1$ per Eq. (2), denoted as $\chi=0,>0,<0$ with different colors. The limit $\Gamma\ll\Delta$ is the “classical” Josephson regime, which is studied in more detail in Supplementary Materials B. Regions with non-zero Chern numbers (hashed) only appear in the regime in which Cooper-pair splitting is active.
@@
that we refer to as chiralities. They are proportionally related to the Josephson current, $I_{J}=I_{c}\sqrt{3}/2\,\chi$, flowing along the kagome wire network, and measure the superconducting phase differences between sublattices. Josephson currents do not flow for $\chi_{ab}=0$, while currents flow maximally clockwise or counter-clockwise when $\chi_{12}=\chi_{23}=\chi_{31}=\mp 1$, respectively. In Fig. 3, we show $\chi\equiv\chi_{12}$ as function of the flux $\phi$ and the ratio $\Gamma/\Delta$, which displays a repeating pattern of values $\chi=0,\pm 1$.
The regime $\Delta\gg\Gamma$ corresponds to the limit where Cooper pair splitting is suppressed [Fig. 1(c) and (d)]; in this case, the superconducting phase relations conform to those expected from Josephson junctions with $E_{J}\sim\Gamma^{2}/\Delta$, as we discuss in detail in the Supplementary Materials B. There we also discuss the emergence of periodicity $6\pi$ (or $3\,\Phi_{0}$) in the dependence in $\phi$ present in Fig. 3 (instead of periodicity $2\pi$).
We next focus on the Chern numbers associated with the filled negative energy single-particle states of Hamiltonian Eq. (1). Fig. 3 shows these Chern numbers, ${\cal C}$, as function of the flux $\phi$ and ratio $\Gamma/\Delta$, indicating trivial and topological phases. The regions in the phase diagram where ${\cal C}=\pm 2$ coincide with those for which the chirality $\chi$ departs from the classical Josephson-array behavior, i.e., those regions where Cooper pair splitting is effective, for $\Gamma$ dominating over $\Delta$. In fact, the dome-like regions have their tip at the “magic” values $\phi=\pm\pi/2,\pm 3\pi/2,\pm 5\pi/2$, flux values for which Cooper pairs interfere destructively going around the triangles of the lattice, since they accrue phases $2\phi=\pm\pi,\pm 3\pi,\pm 5\pi$. In contrast, the fermions (fractions of the pair) do not destructively interfere; such values of flux maximally break time-reversal symmetry.
@@
where $\sigma_{\rm x}$ is a $2\times 2$ Pauli matrix and $\mathds{1}_{p\times p}$ corresponds to a $p\times p$ identity matrix. The two terms in Eq. (3) commute and can be diagonalized separately. The spectrum of $h_{\mathbf{k}}$ is symmetric with respective to zero energy, and thus one of its three eigenvalues is zero for all values of momentum ${\mathbf{k}}$, i.e., it contains a flat band (the model without pairing was studied in Ref. [36]). The other two eigenvalues of $h_{\mathbf{k}}$ are given by ±$\pm\varepsilon_{\mathbf{k}}$ with
@@
±$\pm\Delta\,,\quad-\varepsilon_{\mathbf{k}}\pm\Delta\,,\quad\varepsilon_{\mathbf{k}}\pm\Delta\,.$ (5)
The splitting of bands is shown graphically in Fig. 4. There is a transition point, for $\phi=3\pi/2$ and $\Delta/\Gamma=\sqrt{3}/{2}$, where the middle flat bands at ±$\pm\Delta$ touch the dispersive bands. For $\Delta/\Gamma<\sqrt{3}/{2}$ the Chern number for the filled negative energy bands equals ${\cal C}=-2$, a regime where the system realizes a chiral topological superconductor. Again we remark that Cooper-pair splitting is maximally enhanced for $\Delta\ll\Gamma$ and $\phi=\pm\pi/2,\pm 3\pi/2,\pm 5\pi/2$, where destructive interference suppresses Cooper pair motion around the kagome triangles. When $\Delta\gg\Gamma$ the Chern numbers at half filling are trivial and no topological phase appears. In this limit, the spectrum corresponds to two well-separated copies of the three bands of $h_{\mathbf{k}}$. The total Chern number at half-filling is the sum of those of $h_{\mathbf{k}}$, which sum to zero.
Figure 4: Bands in the flux configuration $\phi=3\pi/2$ for (a) $\Delta=0$, where each one of the three bands is double (spin) degenerate; (b) $20<\Delta<\sqrt{3}\,\Gamma/$ 2, where the degeneracy in each one of the bands is split by ±$\pm\Delta$ as in Eq. (5); and (c) $\Delta>\sqrt{3}\Gamma/2$, where the bands are separated into two sets of spectrum of $h_{\mathbf{k}}$. The light blue and pink colored bands in (b) and (c) indicate anti-symmetric and symmetric spin states, which are non- degenerate when $\Delta\neq 0$.
@@
The Supplementary Material A contains a detailed study of the tight-binding model describing the wire network. Here we summarize the main results. In Fig. 6 we show the spectrum of the wire network for the case in which superconductivity is turned off, for fluxes $\phi=(n-1/2)\pi$, with $n\in\mathbb{Z}$. We observe that at half-filling the spectrum of the three bands around zero energy coincides with that of the effective model, with the flat band at zero energy. The Chern numbers of the band below and above the flat band coincide with those of the effective model (and the Chern numbers of the other core filled levels add to zero). Upon turning on superconductivity, the spectrum assumes the form of that in Eq. (5). Therefore, the more detailed treatment of the wire network yields the same conclusions as the simpler effective model with fewer bands. We stress that this conclusion requires that the spacing between the clusters of bands in Fig. 6 are larger than the superconducting gap $\Delta$. This condition is equivalent to requiring that the length scale of the X-shaped wire crosses is smaller than the superconducting coherence length $\xi$. Recall that we already stated and physically motivated such a condition earlier, when first presenting the effective model.
@@
Figure 6: Band structure of the wire network model for $\Delta=0$, $\gamma=t/2$ and “magic” flux $\phi=\pi/2$. This is a numerical result for $L=4$ along the high-symmetry points $\Gamma-K-M$ in the Brillouin zone. Note the existence of 3x (highlighted in red) degenerate flat bands that sandwich clusters of bands, and are separated from the clusters by a non-zero gap. (These bands are flat for all values of $\phi$.) The inset shows two types of clusters of bands, depicted in blue and red, with the band Chern numbers (±$\pm 1,0$) labeled next to each band. The bands in the middle of both blue and red clusters (depicted in full line in the inset), are flat at the magic values of the flux $\phi=(n-1/2)\pi$, with $n\in\mathbb{Z}$. The non-flat bands depicted in dashed lines are the only ones that carry non-trivial Chern-numbers ±$\pm 1$.
Experimental considerations: Experimentally, it is most natural to impose a uniform magnetic field on the entire lattice. The condition of zero flux per unit cell of the lattice can still be enforced by taking advantage of the $2\pi$-periodic flux dependence. For the kagome configuration, fluxes threading triangles and hexagons must be equal to, respectively, $\phi+2\pi\,m$ and $-2\phi+2\pi\,n$, with $m,n$ integers. For a uniform field, the ratio of these fluxes equals the ratio of the areas $A_{\triangleright}/A_{\varhexagon}=1/6$, yielding discrete values of uniform fields that satisfy the zero-total-flux condition
@@ Acknowledgements.
Supplementary Material A Tight-binding model for wire network

### v2

@@ Abstract
[…]
Introduction: The suppression of dissipation in superconductors is key to building electrical circuits that behave quantum mechanically, making these circuits leading contenders for quantum computing architectures and topological phases [1, 2, 3, 4, 5, 6]. To date, the collection of non-dissipative superconducting circuit elements includes capacitors, inductors, and Josephson junctions. In describing these circuits theoretically, the underlying pairing of electrons is concealed; aside from occasional factors of two in formulas, electron pairing is assumed then forgotten, except insofar as the separation of pairs is regarded as an undesirable source of decoherence and parity non-conservation.
@@
Figure 1: The Y-splitter, the superconducting network component introduced in this work, in two sizes, small (a,c) and large (b,d) compared to the coherence length, $\xi$, and two Josephson coupling strengths, strong (a,b) and weak (c,d). Blue circle represents the coherence length, red segments the insulating barrier thickness. Dominant contribution of Cooper pair transport from the left to the upper arm shown as orange trajectories. The flux $\phi$ through the triangular loop controls interference of single electrons and Cooper pairs. Applying 1/2 (mod 1) flux quanta through the triangular loop leads to destructive interference of Cooper pairs but not single electrons.
@@
The single Y-splitter: The properties of the individual Y-splitter depends on the size of the loop relative to the superconducting coherence length $\xi$, as well as the Josephson coupling strength of its three junctions (assumed to be equal). Four cases are represented in Fig. 1, small (a, c) versus large (b, d) loops, and small (c, d) versus large (a, b) Josephson couplings. In case (a), where the loop is small compared to $\xi$ and couplings are strong (indicated by thin red barriers), Cooper pairs split then recombine while retaining phase coherence, with the phase difference of the separated electrons, accumulated at the junctions, controlled by flux, $\phi$, through the loop, periodic in the single-electron flux quantum, $hc/e$. Cooper pairs can also travel together along each arm, obeying the usual Josephson relations at the junctions within each arm. In case (b), with a large loop compared to $\xi$, coherent pair splitting and recombining is exponentially suppressed. In case (c), coherent splitting is allowed by the small loop but is suppressed compared to pair transport by the weak junctions [thicker red junctions in (c,d)]. This is because transport along a single edge is a second order tunneling process, while transferring two electrons along different paths is a third order process. For case (d), the large loop and weak tunneling doubly suppress coherent splitting. These qualitative descriptions are supported by numerical evidence, as extensively analyzed in Appendix A. There, we examine the behavior of the non-local conductance of a single Y-splitter as a function of the hopping and superconducting parameters. Our results reveal a natural separation into two regimes, which we interpret as indicative of the presence or absence of Cooper-pair splitting. These regimes are in accordance with the limit cases present in Fig. 1.
[…]
In the kagome configuration [Fig. 2(b)], the unit cell area $A_{\rm cell}=A_{\triangleright}+A_{\triangleleft}+A_{\varhexagon}$ comprises two triangles of area $A_{\triangleright}=A_{\triangleleft}$ each and an hexagon of area $A_{\varhexagon}=6\,A_{\triangleright}$. We consider applied perpendicular magnetic fields such that the flux per unit cell is an integer multiple of the flux quantum $\Phi_{0}=\frac{hc}{e}$, a point to which we return when discussing the experimental considerations for such systems. In this situation, we can study model Hamiltonians that do not demand the use of magnetic super cells (in contrast to Hofstadter-like models. This case of zero net flux (mod one flux quantum) per unit cell is equivalent to flux $\phi$ per triangle and $2\phi$ per hexagon. (We henceforth work in units in which the flux $\phi$ is an angular variable, $\phi=2\pi\Phi/\Phi_{0}$, where $\Phi$ is the physical flux through the triangles, pointing into the page.)
@@
Figure 3: Phase diagram of the effective model. The diagram shows the chirality $\chi$ and Chern numbers $\cal{C}$ as functions of the flux $\phi$ and ratio $\Gamma/\Delta$. The chirality takes specific values $\chi_{23}=0,\pm 1$ per Eq. (2), denoted as $\chi=0,>0,<0$ with different colors. The limit $\Gamma\ll\Delta$ is the “classical” Josephson regime, which is studied in more detail in Appendix C. Regions with non-zero Chern numbers (hashed) only appear in the regime in which Cooper-pair splitting is active.
@@
that we refer to as chiralities. They are proportionally related to the Josephson current, $I_{J}=I_{c}\sqrt{3}/2\,\chi$, flowing along the kagome wire network, and measure the superconducting phase differences between sublattices. Josephson currents do not flow for $\chi_{ab}=0$, while currents flow maximally clockwise or counter-clockwise when $\chi_{12}=\chi_{23}=\chi_{31}=\mp 1$, respectively. These equalities suggest that $\mathbb{Z}_{3}$ rotation symmetry is preserved in the low-energy physics of the system. In Fig. 3, we show $\chi\equiv\chi_{12}$ as function of the flux $\phi$ and the ratio $\Gamma/\Delta$, which displays a repeating pattern of values $\chi=0,\pm 1$.
The regime $\Delta\gg\Gamma$ corresponds to the limit where Cooper pair splitting is suppressed [Fig. 1(c) and (d)]; in this case, the superconducting phase relations conform to those expected from Josephson junctions with $E_{J}\sim\Gamma^{2}/\Delta$, as we discuss in detail in the Appendix C. There we also discuss the emergence of periodicity $6\pi$ (or $3\,\Phi_{0}$) in the dependence in $\phi$ present in Fig. 3 (instead of periodicity $2\pi$).
We next focus on the Chern numbers associated with the filled negative energy single-particle states of Hamiltonian Eq. (1). The Chern numbers count the number of monopoles in momentum space present in the many-body ground state, resulting in the quantized topological invariants
$\mathcal{C}=\frac{1}{2\pi}\sum_{\text{filled bands}}\int_{\operatorname{BZ}}d^{2}\mathbf{k}\,\left[\partial_{k_{x}}A_{y}-\partial_{k_{y}}A_{x}\right],$ (3)
where the integral is performed in the first Brillouin zone $\operatorname{BZ}$, and $A_{i}=\langle\,n(\mathbf{k})\,|\partial_{k_{i}}|\,n(\mathbf{k})\,\rangle$ is the Berry connection. In the presence of boundaries, the sign of $\mathcal{C}$ determines the direction of the edge currents: they circulate counterclockwise for $\mathcal{C}>0$ and clockwise for $\mathcal{C}<0$. Fig. 3 shows these Chern numbers $\cal C$}, as function of the flux $\phi$ and ratio $\Gamma/\Delta$, indicating trivial and topological phases. The regions in the phase diagram where ${\cal C}=\pm 2$ coincide with those for which the chirality $\chi$ departs from the classical Josephson-array behavior, i.e., those regions where Cooper pair splitting is effective, for $\Gamma$ dominating over $\Delta$. In fact, the dome-like regions have their tip at the “magic” values $\phi=\pm\pi/2,\pm 3\pi/2,\pm 5\pi/2$, flux values for which Cooper pairs interfere destructively going around the triangles of the lattice, since they accrue phases $2\phi=\pm\pi,\pm 3\pi,\pm 5\pi$. In contrast, the fermions (fractions of the pair) do not destructively interfere; such values of flux maximally break time-reversal symmetry.
@@
where $\sigma_{\rm x}$ is a $2\times 2$ Pauli matrix and $\mathds{1}_{p\times p}$ corresponds to a $p\times p$ identity matrix. The two terms in Eq. (4) commute and can be diagonalized separately. The spectrum of $h_{\mathbf{k}}$ is symmetric with respective to zero energy, and thus one of its three eigenvalues is zero for all values of momentum $\mathbf{k}$}, i.e., it contains a flat band (the model without pairing was studied in Ref. [38]). The other two eigenvalues of $h_{\mathbf{k}}$ are given by $\pm\varepsilon_{\mathbf{k}}$ with
@@
$\pm\Delta\,,\quad-\varepsilon_{\mathbf{k}}\pm\Delta\,,\quad\varepsilon_{\mathbf{k}}\pm\Delta\,.$ (6)
The splitting of bands is shown graphically in Fig. 4. There is a transition point, for $\phi=3\pi/2$ and $\Delta/\Gamma=\sqrt{3}/{2}$, where the middle flat bands at $\pm\Delta$ touch the dispersive bands. For $\Delta/\Gamma<\sqrt{3}/{2}$ the Chern number for the filled negative energy bands equals ${\cal C}=-2$, a regime where the system realizes a chiral topological superconductor. Again we remark that Cooper-pair splitting is maximally enhanced for $\Delta\ll\Gamma$ and $\phi=\pm\pi/2,\pm 3\pi/2,\pm 5\pi/2$, where destructive interference suppresses Cooper pair motion around the kagome triangles. When $\Delta\gg\Gamma$ the Chern numbers at half filling are trivial and no topological phase appears. In this limit, the spectrum corresponds to two well-separated copies of the three bands of $h_{\mathbf{k}}$. The total Chern number at half-filling is the sum of those of $h_{\mathbf{k}}$, which sum to zero.
Figure 4: Bands in the flux configuration $\phi=3\pi/2$ for (a) $\Delta=0$, where each one of the three bands is double (spin) degenerate; (b) $0<\Delta<\sqrt{3}\,\Gamma/2$, where the degeneracy in each one of the bands is split by $\pm\Delta$ as in Eq. (6); and (c) $\Delta>\sqrt{3}\Gamma/2$, where the bands are separated into two sets of spectrum of $h_{\mathbf{k}}$. The light blue and pink colored bands in (b) and (c) indicate anti-symmetric and symmetric spin states, which are non- degenerate when $\Delta\neq 0$.
@@
The Appendix B contains a detailed study of the tight-binding model describing the wire network. Here we summarize the main results. In Fig. 6 we show the spectrum of the wire network for the case in which superconductivity is turned off, for fluxes $\phi=(n-1/2)\pi$, with $n\in\mathbb{Z}$. We observe that at half-filling the spectrum of the three bands around zero energy coincides with that of the effective model, with the flat band at zero energy. The Chern numbers of the band below and above the flat band coincide with those of the effective model (and the Chern numbers of the other core filled levels add to zero). Upon turning on superconductivity, the spectrum assumes the form of that in Eq. (6). Therefore, the more detailed treatment of the wire network yields the same conclusions as the simpler effective model with fewer bands. We stress that this conclusion requires that the spacing between the clusters of bands in Fig. 6 are larger than the superconducting gap $\Delta$. This condition is equivalent to requiring that the length scale of the X-shaped wire crosses is smaller than the superconducting coherence length $\xi$. Recall that we already stated and physically motivated such a condition earlier, when first presenting the effective model.
@@
Figure 6: Band structure of the wire network model for $\Delta=0$, $\gamma=t/2$ and “magic” flux $\phi=\pi/2$. This is a numerical result for $L=4$ along the high-symmetry points $\Gamma-K-M$ in the Brillouin zone. Note the existence of 3x (highlighted in red) degenerate flat bands that sandwich clusters of bands, and are separated from the clusters by a non-zero gap. (These bands are flat for all values of $\phi$.) The inset shows two types of clusters of bands, depicted in blue and red, with the band Chern numbers ($\pm 1,0$) labeled next to each band. The bands in the middle of both blue and red clusters (depicted in full line in the inset), are flat at the magic values of the flux $\phi=(n-1/2)\pi$, with $n\in\mathbb{Z}$. The non-flat bands depicted in dashed lines are the only ones that carry non-trivial Chern-numbers $\pm 1$.
Experimental considerations: Experimentally, it is most natural to impose a uniform magnetic field on the entire lattice. The condition of zero flux per unit cell of the lattice can still be enforced by taking advantage of the $2\pi$-periodic flux dependence. For the kagome configuration, fluxes threading triangles and hexagons must be equal to, respectively, $\phi+2\pi\,m$ and $2\phi+2\pi\,n$, with $m,n$ integers. For a uniform field, the ratio of these fluxes equals the ratio of the areas $A_{\triangleright}/A_{\varhexagon}=1/6$, yielding discrete values of uniform fields that satisfy the zero-total-flux condition
@@ Acknowledgements.
Appendix A Transport properties of a single Y-splitter
In this section we investigate the physics of a single Y-splitter using a tight-binding model for superconducting wires, implemented numerically in Kwant, an open-source Python package [55]. The results offer quantitative insight into the physics of Y-splitters, which is further enhanced through a link between non-local conductance matrix elements and Cooper-pair splitting.
The system is defined on a two-dimensional triangular lattice with lattice spacing $a$, inscribed within an equilateral triangle of edge length $L$ and width $W$, as illustrated in Fig. 7. Three infinitely long normal leads, each of width $W$, are attached at the triangle’s vertices, preserving $\mathbb{Z}_{3}$ rotational symmetry. Each edge of the triangle is separated into two halves that are connected by Josephson barriers (red lines in Fig. 7). Electrons hop across the junctions with amplitude $\gamma$, acquiring a Peierls phase $e^{i\phi/3}$ when moving counterclockwise. Within the same region (other than the barrier) they hop with strength $t$, while an on-site superconducting potential $\Delta$ pairs spin-up and spin-down electrons.
Figure 7: Triangular lattice modeling of a single Y-splitter with flux $\phi$ attached to three leads. Red lines separate two wires and indicate Josephson junctions.
We expect to observe Andreev reflection at the normal-superconductor (NS) interface, where an incident electron reflects as a hole. In a multi-terminal setup, crossed Andreev reflection (CAR) can also occur, in which the reflected hole emerges from a different lead than the one through which the electron entered, requiring the measurement of non-local conductance components. As noted in previous studies [56, 57], CAR can play a crucial role in detecting Cooper-pair splitting.
Let $G_{ij}$ denote the conductance associated with the transmission probability $T_{ij}$ that an electron injected the $i$-th lead travels to the $j$-th one
$G_{ij}=\frac{e^{2}}{h}T_{ij}.$ (8)
The diagonal elements of the conductance matrix, $G_{ii}$, can be determined through local measurements. More relevant to our study are the non-local conductance elements $G_{ij}$ with $i\neq j$, which serve as indicators of Cooper-pair splitting by probing non-local correlations in scattering processes. These elements are also closely linked to crossed Andreev reflection (CAR). Specifically, we define:
$G_{A}\equiv\frac{1}{3}\epsilon_{ij}G_{ij},$
where $\epsilon_{12}=\epsilon_{23}=\epsilon_{31}=1$, $\epsilon_{21}=\epsilon_{32}=\epsilon_{13}=-1$, and zero otherwise. The conductance $G_{A}$ quantifies transport asymmetry, favoring a specific spatial direction dictated by the magnetic flux. We expect $G_{A}$ as a function of $\phi$, $\Delta$, and $\gamma$ to serve as an indicator of the dominant charge carriers in different regimes: for values of $\phi$ that are odd-multiples of $\pi/2$ and sufficiently large correlation lengths, transport is dominated by single electrons, whereas in other cases by Cooper pairs.
We define the dimensionless ratios
${\gamma}_{r}=\gamma/t,\quad{\Delta}_{r}=\Delta/t,\quad{L}_{r}=L/a,$
where energy and length scales are expressed in units of $t$ and $a$, respectively. This formulation is advantageous as it allows us to focus solely on the dependence of $G_{A}$ on the three parameters, together with magetic flux. Fig. 8 presents $G_{A}$ (in units of $e^{2}/h$) as a function of the magnetic flux $\phi$ and the superconducting pairing strength ${\Delta}_{r}$, for a fixed system size ${L}_{r}$ and coupling parameter ${\gamma}_{r}$.
Figure 8: Heatmap of $G_{A}$ as a function of the flux $\phi$ and the ratio $\Delta_{r}$}.
The magnitude of ${\Delta}_{r}$ suggests the existence of two distinct regimes in the system, with $G_{A}$ serving as a non-local order parameter: it takes nontrivial values for ${\Delta}_{r}\ll 1$ and vanishes for ${\Delta}_{r}\gg 1$. The system’s behavior in these two regimes is qualitatively different and can be understood in their respective asymptotic limits. For large ${\Delta}_{r}$, the system behaves as a conventional $s$-wave superconductor carrying supercurrent, whose conductance does not display an asymmetry that is captured by $G_{A}$. In contrast, for small ${\Delta}_{r}$, the physics remains adiabatically connected to that of free electrons, where conductance asymmetry reaches its maximum [36]. In the phase diagram, the nontrivial phase occupies a lobe-shaped region at small ${\Delta}_{r}$, occurring along the lines $\phi=\pm 3\pi/2$. Notably, the flux values where $G_{A}$ reaches its maximum correspond precisely to the half-integer values $\phi=(2n+1)\pi/2$, for $n\in\mathbb{Z}$, where Cooper-pair splitting dominates over paired Cooper-pair transport.
In the following, our goal is to extract the dependence of $G_{A}$ on ${\Delta}_{r}$, ${\gamma}_{r}$, and ${L}_{r}$, providing insight into the system’s behavior. Here, we focus on the line $\phi=3\pi/2$, where Cooper-pair splitting is most prominent. The behavior of $G_{A}$ for fixed ${\gamma}_{r}$, with varying ${L}_{r}$, and vice versa, is shown in Fig. 9. We observe a decay from a large initial value of $G_{A}$ (close to -1) toward zero, forming a peak at $\Delta^{*}_{r}$ that marks the transition out of the Cooper-pair splitting regime. The behavior of $G_{A}$ differs on either side of this point: for ${\Delta}_{r}\gg{\Delta}^{*}_{r}$, it follows an exponential decay, whereas for smaller ${\Delta}_{r}$, it exhibits an algebraic decay. Fig. 10 shows the location of the peak ${\Delta}^{*}_{r}$ for different system sizes ${L}_{r}$ and couplings ${\gamma}_{r}$. In a log-log plot, the dependence of ${\Delta}^{*}_{r}$ is well approximated by a power law. A linear fit to the data yields the relation
${\Delta}^{*}_{r}\sim{L}_{r}^{\alpha}{\gamma}_{r}^{\beta},$
with numerically determined exponents $\alpha=-0.678$ and $\beta=1.223$.
Figure 9: Behavior of $G_{A}$ for fixed (left) system size $L_{r}$ and (right) coupling $\gamma_{r}$ at fixed flux value $\phi=3\pi/2$.
Figure 10: Peak locations of conductance in $\Delta_{r}$-space for varying parameters: (left) different system sizes and (right) different coupling strengths. In a log-log scale, the linear fit indicates a power law relation $\Delta_{r}\sim L_{r}^{\alpha}\gamma_{r}^{\beta}$.
The left panel of Fig. 11 shows the collapse of $G_{A}$ in the early regime when plotted as a function of
$x\equiv{\Delta_{r}}/({L_{r}}^{\alpha}{\gamma_{r}}^{\beta}).$
We observe that the peak occurs approximately at $x\approx 2$. Additionally, the data suggest that $G_{A}$ is largely insensitive to the system size ${L}_{r}$. This is further supported by the right panel, which displays the values of $G_{A}$ for ${\Delta}_{r}=0$, which are well fitted by the function
$g(\gamma_{r})\equiv\frac{G_{A}({\gamma_{r}},\Delta_{r}=0,L_{r})}{e^{2}/h}=\frac{{\gamma_{r}}^{3}}{(1+3{\gamma_{r}}^{2})^{2}}.$
This expression was analytically derived for the case ${L}_{r}=1$ in Appendix A of Ref. [36].
Figure 11: (Left) $G_{A}$ for $\phi=3\pi/2$ as a function of $x\equiv{\Delta_{r}}/{L_{r}}^{\alpha}{\gamma_{r}}^{\beta}$ for various combinations of $L_{r}$} and $\gamma_{r}$}. We observe (i) the collapse of the $G_{A}$ peak around $x\approx 2$ and (ii) that $G_{A}({\Delta_{r}}=0)$ is largely insensitive to ${L}_{r}$ and primarily depends on $\gamma_{r}$}. (Right) The behavior of $G_{A}$ at ${\Delta_{r}}=0$ is well fitted by the analytically derived function $g({\gamma_{r}})$.
Finally, Fig. 12 reveals the exponential decay of $G_{A}$ at large $\Delta_{r}$}. The exponential tail depends exclusively on the product ${\Delta_{r}}\;{L_{r}}$ and is independent of $\gamma_{r}$}. This allows us to capture the general dependence of $G_{A}$ along the line $\phi=3\pi/2$ with the expression:
$\frac{G_{A}({\gamma_{r}},{\Delta_{r}},{L_{r}})}{e^{2}/h}=\frac{16\,{\gamma_{r}}^{3}}{(1+3{\gamma_{r}}^{2})^{2}}f\left(\frac{{\Delta_{r}}}{{L_{r}}^{\alpha}{\gamma_{r}}^{\beta}}\right)\exp(-{\Delta_{r}}\,{L_{r}}),$ (9)
where $\Delta_{r}$}, through the exponential term, introduces a correlation length $\xi\equiv{a}\;{\Delta_{r}}^{-1}$ into the system.
The quantitative dependence of $G_{A}$ on the system parameters is consistent with the qualitative description provided in Fig. 1 of the manuscript. In the limit ${L_{r}}\gg{\Delta_{r}}^{-1}$, as shown in Fig. 1(b), the exponential decay dominates, causing $G_{A}$ to approach zero – reflecting the absence of pair splitting. Conversely, in the limit depicted in Fig. 1(c), where ${\gamma_{r}}^{\beta}\ll{\Delta_{r}}\,{L_{r}}^{\alpha}$, the argument $x\gg 2$ of the function $f(x)$ becomes too large, placing the system outside the relevant pair-splitting regime. This observation aligns with perturbative results, where pair splitting is a third-order process and is suppressed by the dominant second-order process in which electrons travel together.
Figure 12: Exponential decay as a function of $\Delta_{r}\,L_{r}$ for different (left) system sizes $L_{r}$ and (right) hopping coupling $\gamma_{r}$.
Appendix B Tight-binding model for wire network